Sol Summary – Nov 11th

Crew 215 Sol Summary Report 11-NOV-2019
Sol: 1
Summary Title: Operational
Author’s name: Andrew Wheeler

Mission Status: Active

Sol Activity Summary: First sol on Mars saw us hitting the ground running. After a morning brief, we headed south of the hab in search of locations for setting up a sampling grid for micrometeorites and a source of hydrated minerals. No locations for hydrated minerals were identified, however two sites MM5 (518462E 4248949N) and MM6
(519990E 4248036N) for a sampling grid were identified and a grid established at MM5.

This being Remembrance or Armistice Day, activity was interrupted at 1100 with one minute’s silence for the fallen. Lest We Forget. Back at the Hab, the afternoon was spent preparing for fully deploying the sampling grid, troubleshooting space suits and preliminary monitoring of psycho/social stressors in isolated environments. A curry dinner followed by fruit salad rounded out our day.

Look Ahead Plan: Science EVA to MM5 site to set out of the micrometeorite sampling
grid. Monitoring of psycho/social stressors.

Anomalies in work: N/A

Weather: Marginally below freezing overnight, slight breeze, clear skies.

Crew Physical Status: Healthy

EVA: Reconnaissance EVA. See EVA report

Reports to be filed: Sol Summary Report, Operations Report, Science Report, GreenHab Report, EVA Request, EVA Report, Journalist Report, Daily Photos.

Support Requested: See Operations Report

Science Report – Nov 11th

Science Report 11 November 2019
SOL: 1
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Dome Operations: Activity in the science dome included preparing for the deployment of the micrometeorite sampling grid and labeling the sample bags.

Green Hab: Green Hab operations have continued. See Green Hab report.

RAM Operations: The RAM Dome has been activated and has been accessed for suit batteries and fuses for Suit 10 and Suit 2. See the Operations report.

EVA: The first EVA for science involved a geological reconnaissance to the southwest of the hab where two locations on wash areas were scouted for the continuing sampling for micrometeorites. These sites – MM5 (518462E 4248949N) and MM6 (519990E 4248036N) appear to be seasonal alluvial fan deposits and the sampling grid has been established at MM5. See EVA Report.

EVA Suit Maintenance: The crew engineer has evaluated Suit 10 and Suit 2 and has identified the likely faults. See Operations Report.
Additional activities: Nil.

Greenhab Report – Nov 11th

Crew 215 GreenHab Report 11-11-2019
GreenHab Officer: Guy Murphy

Environmental control: Door left open for ventilation 12 midday then reclosed 5 pm.
Average temperatures: –
Hours of supplemental light: Automatic system
Daily water usage for crops: 32 litres (4 full blue watering cans).
Daily water usage for research and/or other purposes:
Water in Blue Tank _____ gallons
Time(s) of watering for crops: 10am, 2pm.
Changes to crops: None
Established plants introduced to Greenhab 03/11/2019

[3x] strawberry plants (Everbearing), in medium metal planter
[2x] small Aloe Vera, in medium metal planter, frost damaged
[1x] spicy orange thyme (thymus fragantissimus), small terra cotta pot
[1x] lemon balm scented geranium (pelargonium x Melissinum), small terra cotta pot
[1x] Mint Mentha – Berries & Cream, small terra cotta pot
[1x] Pineapple Mint – mentha suaveolens variegated, small terra cotta pot
[1x] Ice plant – Delospermo ‘jewel of the desert garnet’, small terra cotta pot
[1x] Sempervium textorum assorted Hen & Chicks, small terra cotta pot
[1x] Rosemary ‘Barbecue’ (Rosmarinus officinalis ‘Barbecue’), large plastic pot
[1x] Rosemary ‘Spice Island’ Rosmarinus officinalis ‘Spice Island’, large plastic pot
[1x] mint, grapefruit, large plastic pot

Seeds Planted in Greenhab 04/11/2019 (Large metal planting tub)

Greek Oregano – Origanum heracleoticum
Chives – Allium schoenoprasum
Sweet Marjoram – Origanum majorana
Lemon Balm – Melissa officialnis
Sage – Savia Officinalis
Thai Basil – Siam Queen
German/Winter Thyme – Thymus vulgaris

Seeds Planted in Greenhab 10/11/2019 (Large metal planting tub)

Coriander – Coriandrum Sativum

Other seeds stored in the Greenhab but not yet planted are identified as –

Onion (Evergreen White Bunching)
Rocket ‘Arugula Rocket Salad ‘Rocky’ (Diplotaxis tenuifolia)
Spinach ‘Lakeside F1’
Tomato Pole Roma ‘San Marzano’ (Solanum lycopersicum)
Carrots ‘Purple Sun F1’
Radish ‘icicle’
Snow Pea (Oregon Giant)
Cucumbers (Patio Snackers F1)
Pepper (Mini Bell Color Mix)
Tomato (Chocolate Cherry)
Perpetual Spinach (Green Leaf Chard)

Harvest: None

Support/supplies needed: N/A

Operations Report – Nov 11th

Crew 215 Operations Report 11-NOV-2019
SOL: 1
Name of person filing report: Shane Usher

Non-nominal systems: Opportunity, Suit 10, Suit 5 Helmet, No power to multiple downstairs wall sockets.

Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning (fuse type 5 is blown), Suit 5 Helmet falling apart, Suit 2 Battery supplying low power. Loft Tank Feed tank leaky (temporarily fixed with gaffer tape).
The water pump is very loud.

Generator: run
Hours run: 12h 30 min
From what time last night: 1800
To what time this morning: 0730
List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before generator is run at night): 78%
Diesel Reading – 54%
Station Propane Reading – 74%
Ethanol Free Gasoline: N/A

Water (loft tank): 55 gallons
Water Meter: 1457225 units
Water (static tank): 304 gallons (not filled)
Static to Loft Pump used – yes
Water in Green Hab: 225 gallons – 28 L (7 gallons) used
Water in Science Dome: 0 gallons

Toilet tank emptied: no

Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Spirit rover used: yes
Hours: 110.0
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 100%
Currently charging: yes.

Opportunity rover used: no
Hours: 65.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: no

Curiosity rover used: yes
Hours: 115.7
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes

Notes on rovers: no
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director

HabCar used and why, where? With DIRECTOR
CrewCar used and why, where? At Grand Junction awaiting Crew 216

General notes and comments: N/A

Summary of internet: Functional today, but CAPCOMs was down in the evening on Sol 0.
Summary of suits and radios: Suit 2 needs a replacement battery, Suit 5 Helmet falling
apart, Suit 10 requires type 5 fuse.
Summary of Hab operations: All systems nominal apart from; Stateroom5, Edison screw lightbulb Max 13W required; Fire alarm test button not working in Hab upper level; Fire alarm test button disintegrated in GreenHab; Numerous powerpoints not functional on the ground floor of hab. The water pump is very loud.
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered (28L). The heater automatically came on at 5 pm.
Summary of Science Dome operations: All systems nominal.
Summary of RAM operations: Not used
Summary of any observatory issues: Nominal
Summary of health and safety issues: All crew healthy

Questions, concerns and requests to Mission Support: Static tank topped up to approximately 350 L. Water heater unusable. Opportunity is currently unusable.
We need advice about how we can obtain a replacement battery and fuse (type 5) for Suits. Loft tank feed pipe needs a new hose connection to eliminate the leak. We might be able to improve the current setup with Teflon tape and duct tape.

Supplemental:

Supply Request List:
12-volt lead-acid battery – Suit 2 Hab ground floor
Fuse (Type 5) – Suit 10 Hab ground floor
Fire smoke alarm – Hab upper level
Fire smoke alarm – GreenHab
Water Pump – water tank outside next to Hab?
Water tank feed hose connection – Loft
Gaffer Tape – RAM
Teflon tape – RAM
Electrical Service: Powerpoint not functioning – Hab ground floor near the hot water system
Safety: Static electric shock becoming significant around the Hab

EVA Report – Nov 11th

***ATTN: We understand the need to give exact coordinates for safety reasons, so they are included below, however, we believe the previous issues with having our grid disturbed were due to posting the exact location online. To prevent unwanted visitors from the sampling site, are we able to replace the final 3-4 coordinate numbers with XXXX when posting in public? Thank you.***

Crew 215 EVA Report 11 Nov 2019
EVA # 01
Author: Jennifer Lane

Purpose of EVA: Geological Reconnaissance

Start time: 0955
End time: 1215

Narrative: The first EVA for Crew 215 saw us scouting for an appropriate location to set up our 10m x 10m grid (MM5) for the micrometeorite sampling, which will be our 2nd such grid since Crew 214 began. We found an undisturbed location at 518462E 4248949N just off Cow Dung Road to begin setting up the grid with the boundaries.

After a minor issue with aligning the strings to NESW (which we attribute to the crumbly Martian regolith and lack of magnetic poles on Mars, rather than human error) we were able to set up the square, ready for the full deployment on subsequent EVAs.

Just after finishing, our alarm went off signaling 11 am, and so we spent a minute in silence for Remembrance or Armistice Day.

We soon continued on our way south and were able to scout another possible location we’re calling MM6. We took the coordinates of this location (519990E 4248036N) and then continued further south along Cow Dung Road before parking at the junction with road 1101 in search of gypsum, which we were unfortunately not able to find in this location, nor was there a site appropriate for any micrometeorite sampling.

Destination: Drove from the hab, south along Cow Dung Road, where we set up our grid, and then further south on Cow Dung Road to road 1101, before returning to the hab.

Coordinates (use UTM NAD27 CONUS): 518462E 4248949N and 519900E 4248030N
Participants: Commander Andrew Wheeler, Larissa Wilson, and Jennifer Lane
Road(s) and routes per MDRS Map: Cow Dung Road to road 1101.
Mode of travel: Spirit Rover and Curiosity Rover

EVA Report – Nov 11th

Cancel this!!!

Corrected EVA report to follow.

On Mon, Nov 11, 2019 at 10:32 PM BMD CapCom <bmdcapcom> wrote:

Mon 11 Nov Sol 1
by Guy Murphy

This morning we woke up on Mars. We emerged from our sleeping quarters at 7 am to views of the Sun rising across an alien landscape through the upper-level Hab windows. For the next 12 days, we are in full simulation mode (sim). We do not go outside of the designated pressurised areas of the campus unless wearing space suits and following a depressurisation protocol. We have to make do with the supplies we have and undertake our research projects and Hab maintenance while not breaking sim.

The crew is following a set daily schedule over the course of the mission, which includes rising at 7 am and lights out at 10 pm. This is to keep everyone working together on the same sleep cycle, and get the most of the available daylight hours. We are fortunate to have relatively warm, clear and dry days forecast for the next fortnight. Typically in this part of Utah, it is much colder at this time of year, with greater odds of precipitation.

The first EVA (Extra-Vehicular Activity) onto the Martian surface was completed this morning, with Larissa and Jennifer taking their first steps on another world. The 11th of November is Remembrance Day in Australia and New Zealand. As is customary, the crew stopped while out in the field for a minute of silence at 11 am to honour Australian and New Zealand service men and women.

For lunch, we consumed the remaining spaghetti bolognese from last night. Early afternoon, I showed the new crew how to make butter, cheese, and bread from the various dry ingredients in our stores. The bread machine is proving its worth.

On a long-duration mission with a limited pantry, we both crave novelty and must avoid wasting food products. Larissa is experimenting with creating a vegetarian mock pulled pork dish derived from banana skins. She has candied some orange and mandarine peel to include in a fruit bread she will bake tomorrow.

EVA Report – Nov 11th

Mon 11 Nov Sol 1
by Guy Murphy

This morning we woke up on Mars. We emerged from our sleeping quarters at 7 am to views of the Sun rising across an alien landscape through the upper-level Hab windows. For the next 12 days, we are in full simulation mode (sim). We do not go outside of the designated pressurised areas of the campus unless wearing space suits and following a depressurisation protocol. We have to make do with the supplies we have and undertake our research projects and Hab maintenance while not breaking sim.

The crew is following a set daily schedule over the course of the mission, which includes rising at 7 am and lights out at 10 pm. This is to keep everyone working together on the same sleep cycle, and get the most of the available daylight hours. We are fortunate to have relatively warm, clear and dry days forecast for the next fortnight. Typically in this part of Utah, it is much colder at this time of year, with greater odds of precipitation.

The first EVA (Extra-Vehicular Activity) onto the Martian surface was completed this morning, with Larissa and Jennifer taking their first steps on another world. The 11th of November is Remembrance Day in Australia and New Zealand. As is customary, the crew stopped while out in the field for a minute of silence at 11 am to honour Australian and New Zealand service men and women.

For lunch, we consumed the remaining spaghetti bolognese from last night. Early afternoon, I showed the new crew how to make butter, cheese, and bread from the various dry ingredients in our stores. The bread machine is proving its worth.

On a long-duration mission with a limited pantry, we both crave novelty and must avoid wasting food products. Larissa is experimenting with creating a vegetarian mock pulled pork dish derived from banana skins. She has candied some orange and mandarine peel to include in a fruit bread she will bake tomorrow.

Journalist Report – Nov 11th

Mon 11 Nov Sol 1
by Guy Murphy

This morning we woke up on Mars. We emerged from our sleeping quarters at 7 am to views of the Sun rising across an alien landscape through the upper-level Hab windows. For the next 12 days, we are in full simulation mode (sim). We do not go outside of the designated pressurised areas of the campus unless wearing space suits and following a depressurisation protocol. We have to make do with the supplies we have, and undertake our research projects and Hab maintenance while not breaking sim.

The crew is following a set daily schedule over the course of the mission, which includes rising at 7 am and lights out at 10 pm. This is to keep everyone working together on the same sleep cycle, and get the most of the available daylight hours.

We are fortunate to have relatively warm, clear and dry days forecast for the next fortnight. Typically in this part of Utah it is much colder at this time of year, with greater odds of precipitation.
The first EVA (Extra-Vehicular Activity) onto the Martian surface was completed this morning, with Larissa and Jennifer taking their first steps on another world.

The 11th of November is Remembrance Day in Australia and New Zealand. As is customary, the crew stopped while out in the field for a minute of silence at 11 am to honour Australian and New Zealand service men and women.

For lunch, we consumed the remaining spaghetti Bolognese from last night. Early afternoon, I showed the new crew how to make butter, cheese, and bread from the various dry ingredients in our stores. The bread machine is proving its worth.

On a long-duration mission with a limited pantry, we both crave novelty and must avoid wasting food products. Larissa is experimenting with creating a vegetarian mock pulled pork dish derived from banana skins. She has candied some orange and mandarine peel to include in a fruit bread she will bake tomorrow.